# Designing Simple Mechanisms

### v1

@@
Shengwu LiI thank Eric Chen for valuable comments.
@@
May 2, 2024)
Which mechanisms are simple to play? When is it easy for participants to see that a mechanism is incentive-compatible? I will start by explaining how and why economists came to ask these questions. Then I will discuss three recent answers, that capture different aspects of what makes a mechanism simple.
@@
Hurwicz’s mechanism is an abstract representation of richer real-world institutions. The dials stand for purchases that the agents could make, bids that they could place, bargaining strategies that they could adopt, and so on. Later work converged on a standard abstraction, in the form of the revelation principle: When studying what might happen in an arbitrary equilibrium of an arbitrary mechanism, it is without loss of generality to study what can be achieved by incentive-compatible revelation mechanisms. In such mechanisms, each agent is asked to directly report their private information, and it is in their best interest to report truthfully (Gibbard, 1973; Dasgupta et al., 1979; Myerson, 1979). A mechanism is Bayesian incentive-compatible if reporting truthfully is a Bayesian Nash equilibrium. A mechanism is dominant-strategy incentive compatible (‘strategy-proof’) if reporting truthfully is always a best response, regardless of the other players’ behavior.
@@
On the metaphorical interpretation, it does not matter whether participants understand that the mechanism is incentive-compatible. After all, people do not interact with the mechanism per se; instead, the mechanism represents aspects of their everyday economic transactions. People may be proficient at arranging those transactions—making purchases in shops, haggling over contracts—and yet not recognize that the mechanism is incentive-compatible when confronted with its abstract form.
The limits of this interpretation became clear when economists began to design the rules for real-world marketplaces. This started in the 1990s, with auctions for radio spectrum broadcast licenses and the National Resident Matching Program (NRMP), a clearinghouse that matches American medical students to their first jobs (Milgrom, 2000; Roth, 2002). The practice soon found myriad applications; for instance, economists designed systems to coordinate electricity generation, to match students to public schools, to swap donated goods between food banks, and to sell loans from the Bank of England to financial institutions (Wilson, 2002; Abdulkadiroğlu et al., 2005; Prendergast, 2022; Klemperer, 2010).
@@
The practice of market design led economists to focus on a more literal interpretation of mechanisms, namely that mechanism design involves “explicitly analyzing the consequences of trading rules that presumably are really common knowledge” (Wilson, 1987). On this view, the rules of the mechanism are not shorthand for a vague decentralized process; rather, they capture the actual rules of the institution as understood by the participants. For instance, the rules for a multi-round auction specify what bids are permitted, what each bidder learns about the bids from past rounds, when the auction will end, and how the winner and the payments will be determined.
@@
When we regard the mechanism as capturing the rules of a formal process, such as an auction or a school choice system, it makes sense to ask whether participants understand that the mechanism is incentive-compatible. This question is particularly pressing for the purposes of market design, for several reasons. First, market designers often consider novel design proposals that are unfamiliar to the intended participants—or worse, may be deceptively familiar, leading participants to adopt heuristics that are unsuited to the new rules. Second, in applications such as school choice, people participate in the mechanism exactly once, and thus cannot learn from experience. Third, even if the designer asserts that the mechanism is incentive-compatible, participants may distrust that claim. And that distrust may be justified: Before the intervention of economists, the NRMP claimed that its algorithm made it incentive-compatible for medical students to report their true preferences, even though this was false (Williams, 1995). [^1] At the time, the NRMP ran the hospital-proposing deferred acceptance algorithm, which is not incentive-compatible for the applicants. Similarly, when Google started selling Internet advertising via the Generalized Second-Price auction, it incorrectly claimed that it was a dominant strategy for participants to bid their value, asserting that its “unique auction model uses Nobel Prize-winning economic theory to eliminate […] that feeling you’ve paid too much” (Edelman et al., 2007). (This presumably referred to the 1996 prize to William Vickrey and James Mirrlees, for work that we will shortly discuss.)
@@
Participants in real-world mechanisms do not always recognize that the mechanism is incentive-compatible. The modern NRMP algorithm is (for all practical purposes) incentive-compatible [^2] The NRMP uses the applicant-proposing deferred acceptance algorithm, modified to account for couples. The standard deferred acceptance algorithm is strategy-proof for the proposing side (Dubins and Freedman, 1981), but accounting for couples slightly undermines this property. Computational experiments indicate that the probability that an applicant can profitably deviate under the NRMP algorithm is about $1$ in $10,000$ (Roth and Peranson, 1999)., and the NRMP prominently advises that participants should rank jobs in their preferred order. Nonetheless, 17% of graduating medical students claim, when surveyed, to have submitted a non-truthful rank-order list to the NRMP (Rees-Jones, 2018). Rees-Jones and Skowronek (2018) recruited medical students who recently participated in the NRMP, to be subjects in an incentived experiment with the same NRMP algorithm. They found that a similar proportion (23%) of subjects submit non-truthful rank-order lists in the experiment. As further evidence, in strategy-proof mechanisms that match students to degree programs, some students make unambiguous mistakes. They rank a program without a scholarship above the same program with a scholarship, even though the scholarship is worth thousands of euros (Hassidim et al., 2021; Shorrer and Sóvágó, 2023).
@@
Consequently, it is often not enough that a mechanism is incentive-compatible; participants have to see for themselves that it is incentive-compatible; the mechanism has to be simple. Simplicity bypasses the need for participants to trust the designer. Simplicity eases the cognitive cost of participation, which is just as real as other costs. Simplicity can level the playing field, in the sense of Pathak and Sönmez (2008), preventing unfair outcomes caused by unequal strategic sophistication. Finally, simplicity allows the designer to rely more confidently on the predictions of classical game theory, because that analysis depends on participants correctly responding to incentives. [^3] Li and Dworczak (2024) make a formal critique of this last justification. If participants are strategically unsophisticated, then instead of using a simple mechanism, the designer might profitably adopt complex mechanisms that confuse the participants.
In the rest of this article, I explain three challenges that participants may need to overcome, in order to recognize incentive compatibility. These are:
  1. 1.
    Thinking contingently about unobserved opponent moves.
  2. 2.
    Planning for their own future moves.
  3. 3.
    Reasoning about other players’ beliefs.
In each of these cases, I will explain formal criteria that capture the difficulty in question, and discuss mechanisms that alleviate that difficulty.
@@
The three challenges above are not exhaustive of the ways that mechanisms can be complex. This article is not a comprehensive survey, and it is shaped by my idiosyncratic tastes and the limits of my expertise. I have omitted excellent papers for the sake of brevity.
1 Thinking contingently about unobserved moves
@@
In his seminal paper, Vickrey (1961) invented [^4] It is more accurate to say that Vickrey reinvented the second-price auction. Second-price auctions were used to sell postage stamps to collectors as early as 1893 (Lucking-Reiley, 2000). the second-price auction: All participants simultaneously place sealed bids, then the highest bidder wins and pays the second-highest bid. Vickrey argued that the second-price auction is desirable because it is simple:
Each bidder can confine his efforts and attention to an appraisal of the value the article would have in his own hands, at a considerable saving in mental strain and possibly in out-of-pocket expense.
@@
The second-price auction is strategy-proof. One can see this by analogy. In an ascending auction, the price starts low and gradually rises. At each moment, each bidder decides whether to keep bidding or to quit irrevocably. When just one bidder is left, that bidder clinches the object at the current price. It is obvious that you should keep bidding if the price is below your value, and should quit if the price is above.
Vickrey pointed out that the second-price auction is “logically isomorphic to” the ascending auction. In the ascending auction, each bidder essentially chooses when to quit. The bidder with the highest quit-price wins, and pays the second-highest quit-price. Quit-prices in the ascending auction are equivalent to bids in the second-price auction, so it is a dominant strategy to place a bid equal to your value.
@@
Notice that Vickrey’s argument runs the revelation principle ‘in reverse’. The second-price auction is an incentive-compatible revelation mechanism. We have reasoned about that mechanism by studying a dynamic mechanism—the ascending auction—and arguing that two mechanisms are equivalent. But do real participants treat them as equivalent?
@@
For auctions in the wild, it can be hard to tell whether participants are bidding ‘truthfully’, because we usually do not know their value for the object. To overcome this obstacle, Kagel et al. (1987) had lab subjects bid in auctions for a ‘virtual prize’, which had a different value for each bidder. Each subject knew their own value, in dollars, for the virtual prize. The winning bidder received cash equal to their value minus the price they paid. This method enables the experimenter to observe whether subjects are bidding truthfully. [^5] Outside of revelation mechanisms, whether a strategy is ‘truthful’ is somewhat a matter of labeling. In such cases, I use ‘truthful’ to refer to the strategy that the designer intends for the participants to adopt.
@@
Kagel et al. (1987) randomized whether subjects participated in ascending auctions or in second-price auctions. In theory, these two formats are equivalent; and it is a dominant strategy to choose a bid (or quit-price) equal to your value. [^6] Moreover, the money lottery induced by bidding your value first-order stochastically dominates the lottery from any other bid. Thus, bidding truthfully maximizes expected utility regardless of your risk preferences. Thus, we might expect that in both formats subjects will bid truthfully, and that the clearing price will always be equal to the second-highest value.
@@
The theoretical prediction was largely borne out in ascending auctions: Subjects rapidly figured out that truthful bidding is optimal, and in 76% of auctions the clearing price was at the second-highest value (plus or minus one increment of the price clock).
@@
By contrast, in second-price auctions, subjects did not behave as theory predicted. Subjects often submitted bids far from their values, the clearing price was at the second-highest value (plus or minus one increment) in no more than 20% of auctions. Subjects played the second-price auction 30 times in a row, with their values drawn randomly each time, but prices remained far from the theoretical prediction.
@@
In summary, Kagel et al. (1987) found that lab subjects rapidly converge on truthful bidding in ascending auctions, but make large and persistent mistakes in second-price auctions. This result is well-replicated. [^7] For replications, see Kagel and Levin (1993), McCabe et al. (1990), Harstad (2000), Li (2017), and Breitmoser and Schweighofer-Kodritsch (2022). Ascending auctions and second-price auctions are not equivalent in practice.
@@
What did our theory miss? What makes it easy for people to see that ascending auctions are strategy-proof? What makes it hard to see this for second-price auctions?
@@
Suppose a participant wants to check directly that truthful bidding weakly dominates some deviating bid in a second-price auction; then they have to compare the payoffs for each profile of opponent bids. To bid optimally, they have to understand that if a truthful bid of \$40 would win at a price of \$30, then deviating to bid \$50 would also lead to a win at \$30. And if a truthful bid of \$40 would lose, then a bid of \$50 could only win at prices above \$40. If the bidder does not carefully keep track of each contingency, then they might be tempted to bid above their value in order to raise their chance of winning. The mistake is that this raises their chance of winning only when the price is so high that winning is undesirable.
@@
By contrast, suppose that a participant is bidding in an ascending auction, and values the object at \$40. The price is now \$30. If the participant continues to bid truthfully, then they will either win at a price between \$30 and \$40, or they will lose and have payoff zero. Thus, every possible payoff from bidding truthfully is at least zero. On the other hand, if they deviate to quit right now, then their payoff is zero for sure. [^8] We have ignored the possibility of ties. Formally, we are modeling the ascending auction as a mechanism in which we cycle between the active bidders in some fixed order, asking each to raise their bid by a dollar or to quit irrevocably. When just one bidder remains, that bidder wins and pays their last bid. The worst-case payoff from truthful bidding is at least as good as the best-case payoff from deviating, so the participant need not keep track of different contingencies to see that truthful bidding is optimal.
@@
A symmetric argument establishes that the participant should quit when the price exceeds their value. Suppose that the participant values the object at \$40 and the price rises to \$41. If they quit, then their payoff is zero for sure. If instead they keep bidding, then their payoff is no more than zero, because now they can only win at prices above \$41.
In hindsight, the distinction between second-price and ascending auctions looks trivial: To play optimally in a second-price auction, you need to compare strategies case-by-case [^9] Unless you notice, as Vickrey did, the isomorphism to the ascending auction. In a lab experiment, Breitmoser and Schweighofer-Kodritsch (2022) found that subjects are more likely to bid truthfully in second-price auctions when the mechanism is framed in a way that makes the isomorphism salient., whereas to play optimally in an ascending auction, you need not make that comparison.
@@
We can generalize this idea to any extensive-form mechanism. Strategy $S_{n}$ obviously dominates strategy $S^{\prime}_{n}$ if at any information set where $S_{n}$ and $S^{\prime}_{n}$ choose different actions for the first time, the worst-case payoff under $S_{n}$ is at least as good as the best-case payoff under $S^{\prime}_{n}$. [^10] This results in a partial ranking of strategies, that can be understood as a weakening of a standard monotonicity axiom for choice under uncertainty (Echenique et al., 2022). (When the worst case from $S_{n}$ is preferred to the best case from $S^{\prime}_{n}$, a case-by-case comparison is not needed.) As we just saw, in ascending auctions, the truthful strategy obviously dominates any other strategy. A mechanism is obviously strategy-proof (OSP) if the truthful strategy is obviously dominant (Li, 2017).
@@
Some strategy-proof mechanisms are not obviously strategy-proof. At the moment you submit your bid in a second-price auction, bidding your value can lead to a zero payoff, whereas bidding above your value can yield a positive payoff. Thus, truthful bidding is not an obviously dominant strategy.
@@
Obvious strategy-proofness pins down the extensive form of the ascending auction. That is, suppose we restrict attention to auctions that are efficient (allocating the object to a bidder with maximal value) and in which only the winning bidder makes payments. In that class, an extensive-form mechanism is OSP if and only if it is an ascending auction (Li and Lo, 2023, p. 576–582). [^11] Alternatively, the ascending auction can be characterized as the only strategy-proof mechanism that is incentive-compatible for the auctioneer (Akbarpour and Li, 2020). Under weaker restrictions, one can characterize OSP mechanisms as ‘personal-clock auctions’, a larger class that includes ascending auctions, descending-price reverse auctions, and some hybrid formats (Li, 2017). Roughly, in a personal-clock auction each participant chooses between clinching a sure thing (e.g. leaving empty-handed) and a tentative alternative with a price that gets steadily worse for the participant (e.g. bidding to buy the object at a price that increases each round).
The 2017 Incentive Auction conducted by the US Federal Communications Commission (FCC) transacted US\$19.8 billion of radio spectrum rights. It included a reverse auction that paid TV stations to relinquish their over-the-air broadcast rights. This design problem involved computationally hard interference constraints, which made it infeasible to run a Vickrey-Clarke-Groves mechanism (Leyton-Brown et al., 2017). Instead, the proposed auction format used a suite of state-of-the-art feasibility-checking algorithms to determine the allocation and payments. Because of this complexity, it was especially important for participants to see that the proposed auction was incentive-compatible.
@@
The reverse auction proceeds, roughly, as follows: The FCC makes an initial offer to each TV station. Each TV station can either quit or continue. If it quits, then it remains on the air, retaining its broadcast rights with zero net payment. If it continues, then either it sells its rights at the current price, or the FCC makes a lower offer and the process repeats. It is an obviously dominant strategy for each TV station to bid truthfully, quitting once the price drops below its value for remaining on the air. This reverse auction is another example of a personal-clock auction. It belongs to the class of deferred-acceptance clock auctions studied by Milgrom and Segal (2020), which have useful implications for privacy and budget-balance.
@@
You can see that the FCC’s reverse auction is incentive-compatible, even though I have not described how the offers are determined or when the auction stops. [^12] These details depend on those complicated feasibility-checking algorithms. Most bidders lacked the computational horsepower to check whether the algorithms were running as intended. This is no accident. A mechanism is OSP if and only if truth-telling can be seen to be dominant from a certain kind of partial description. Such a description specifies the sequences of queries that one participant might receive, the answers that are allowed, and how queries and answers map to possible outcomes (Li, 2017, Theorem 1). Notice that this description omits how the outcome depends on unobserved opponent moves; in this sense, OSP mechanisms formalize the idea of dominance that can be recognized without contingent reasoning.
@@
Psychologically, it is harder to account for hypothetical contingencies, than to reason about observed events (Esponda and Vespa, 2014). Dynamic mechanisms can help participants to avoid mistakes, by resolving uncertainty about what other players have done. This connects to a larger literature on contingent reasoning (Martínez-Marquina et al., 2019; Cohen and Li, 2022; Esponda and Vespa, 2023), reviewed by Niederle and Vespa (2023).

### v3

@@
Shengwu LiI thank Tilman Borgers, Eric Chen, and Jiangtao Li for valuable comments.
@@
July 18, 2024)
Economists today design rules for real-world marketplaces. This began in the 1990s, when economists designed auctions for radio spectrum broadcast licenses and reformed the National Resident Matching Program (NRMP), a clearinghouse that matches graduates of American medical schools to their first jobs (Milgrom, 2000; Roth, 2002). The practice soon found myriad applications; for instance, economists designed systems to coordinate electricity generation, to match students to public schools, to swap donated goods between food banks, and to sell loans from the Bank of England to financial institutions (Wilson, 2002; Abdulkadiroğlu et al., 2005; Prendergast, 2017, 2022; Klemperer, 2010).
In these mechanisms, participants convey information about their preferences, which affects who gets what and (sometimes) how much they pay. Participants will naturally try to figure out whether it makes sense to provide information in accordance with their true preferences, or whether it might prove advantageous to strategize—that is whether the mechanism is incentive-compatible. If a mechanism is simple to play, then it will be easier for participants to see that it is in their interest to reveal their true preferences.
But what makes a mechanism simple? When is it easy for participants to see that a mechanism is incentive-compatible? I will start here by explaining how and why economists came to ask these questions. Then I will discuss three recent answers, which will capture different aspects of what makes a mechanism simple.
1 Taking mechanisms literally
@@
Hurwicz is describing an abstract representation of richer real-world institutions. The dials stand for purchases that the agents could make, bids that they could place, bargaining strategies that they could adopt, and so on. Later work converged on a standard abstraction, in the form of the “revelation principle”: Under some conditions, it is without loss of generality to restrict attention to truthful behavior in incentive-compatible revelation mechanisms. In such mechanisms, each agent is asked to directly report their private information, and it is in their best interest to report truthfully (Gibbard, 1973; Dasgupta et al., 1979; Myerson, 1979). A mechanism is dominant-strategy incentive-compatible (‘strategy-proof’) if reporting truthfully is always a best response, regardless of the other players’ behavior.
@@
On the metaphorical interpretation of a mechanism, it does not matter whether participants understand that the mechanism is incentive-compatible. After all, people do not interact with the mechanism per se; instead, the mechanism represents aspects of their everyday economic transactions. People may be proficient at arranging those transactions—making purchases in shops, haggling over contracts—and yet not recognize that the mechanism is incentive-compatible when confronted with its abstract form.
@@
The practice of designing real-world markets led economists to focus on a more literal interpretation of mechanisms, namely that mechanism design involves “explicitly analyzing the consequences of trading rules that presumably are really common knowledge” (Wilson, 1987). On this view, the rules of the mechanism are not shorthand for a vague decentralized process; rather, they capture actual rules of the institution as understood by participants. For instance, the rules for a sealed-bid auction specify what bids are permitted and how the winner and the payments are determined.
@@
When we regard the mechanism as capturing the rules of a formal process, such as an auction or a school choice system, it matters whether participants understand that the mechanism is incentive-compatible. This is a pressing question for market design, for several reasons. First, market designers often consider novel design proposals that are unfamiliar to the intended participants—or worse, may be deceptively familiar, leading participants to adopt heuristics that are unsuited to the new rules. Second, in applications such as school choice or certain high-stakes auctions, some people participate in the mechanism exactly once, and thus cannot learn from experience. Third, even if the designer asserts that the mechanism is incentive-compatible, participants may distrust that claim. Moreover, that distrust may be justified: before the intervention of economists, the National Resident Matching Program (mistakenly) claimed that its algorithm made it incentive-compatible for medical students to report their true preferences, even though this was false (Williams, 1995). [^1] At the time, the National Resident Matching Program ran the hospital-proposing deferred acceptance algorithm, which is not incentive-compatible for the applicants. Similarly, when Google started selling Internet advertising via auction, it incorrectly claimed that its auction format was strategy-proof, asserting that its “unique auction model uses Nobel Prize-winning economic theory to eliminate [. . . ] that feeling you’ve paid too much” (Edelman et al., 2007). (This presumably referred to the 1996 prize to William Vickrey and James Mirrlees, for work that we will shortly discuss.) To the best of my knowledge, both claims were genuine mistakes rather than intentional deception.
@@
Participants in real-world mechanisms do not always recognize that the mechanism is incentive-compatible. The modern algorithm for the National Resident Matching Program is (for all practical purposes) incentive-compatible [^2] The program now uses the applicant-proposing deferred acceptance algorithm, modified to account for couples. The standard deferred acceptance algorithm is strategy-proof for the proposing side (Dubins and Freedman, 1981), but accounting for couples slightly undermines this property. Computational experiments indicate that the probability that an applicant can profitably deviate under the modern algorithm is about $1$ in $10,000$ (Roth and Peranson, 1999)., and the program prominently advises that applicants should rank jobs in their preferred order. Nonetheless, 17 percent of applicants claimed, when surveyed, to have submitted a non-truthful rank-order list (Rees-Jones, 2018). When Rees-Jones and Skowronek (2018) recruited medical students who recently participated in the National Resident Matching Program to take part in an incentivized experiment with the same algorithm, they found that 23 percent of subjects submitted non-truthful rank-order lists in the experiment. As further evidence, in strategy-proof mechanisms that match students to degree programs, some students make unambiguous mistakes. They rank a program without a scholarship above the same program with a scholarship, even though the scholarship is worth thousands of euros. This behavior has been found in matching mechanisms for graduate psychology degrees in Israel (Hassidim et al., 2021) and for undergraduate degrees in Hungary (Shorrer and Sóvágó, 2023).
@@
Consequently, it is often not enough that a mechanism is theoretically incentive-compatible. Participants have to see for themselves that it is incentive-compatible; the mechanism has to be simple. Simplicity bypasses the need for participants to trust the mechanism designer. Simplicity eases the cognitive cost of participation, which is just as real as other costs. Simplicity can level the playing field, in the sense of Pathak and Sönmez (2008), by preventing unfair outcomes caused by unequal strategic sophistication. Finally, simplicity allows the designer to rely more confidently on the predictions of classical game theory, because that analysis depends on participants responding predictably and correctly to incentives. [^3] Li and Dworczak (2024) make a formal critique of this last justification. If participants are strategically unsophisticated, then instead of using a simple mechanism, the designer might profitably adopt complex mechanisms that confuse the participants.
In the rest of this article, I explain three challenges that participants may need to overcome, in order to recognize incentive compatibility: 1) thinking contingently about unobserved moves by other players; 2) planning for their own future moves; and 3) reasoning about other players’ beliefs. In each of these cases, I will explain formal criteria that capture the difficulty in question, and discuss mechanisms that alleviate that difficulty.
@@
These three challenges are not exhaustive of the ways that mechanisms can be simple or complex. This article is not a comprehensive survey, and it is shaped by my idiosyncratic tastes and the limits of my expertise. I have omitted excellent papers for the sake of brevity.
2 Thinking contingently about unobserved moves by other players
Some mechanisms require participants to reason case-by-case about other players’ moves, in order to see that the mechanism is incentive-compatible.
@@ 2 Thinking contingently about unobserved moves by other players
In a seminal paper, Vickrey (1961) invented [^4] It is more accurate to say that Vickrey reinvented the second-price auction. Second-price auctions were used to sell postage stamps to collectors as early as 1893 (Lucking-Reiley, 2000). the second-price sealed-bid auction, in which all participants simultaneously place bids, then the highest bidder wins and pays the second-highest bid. Vickrey argued that the second-price auction is desirable because it is simple: “Each bidder can confine his efforts and attention to an appraisal of the value the article would have in his own hands, at a considerable saving in mental strain and possibly in out-of-pocket expense.”
@@
The second-price sealed-bid auction is strategy-proof, so it is in your own best interest to bid your true value. One can see this by analogy. In a dynamic ascending auction, the price starts low and gradually rises. At each moment, each bidder decides whether to keep bidding or to quit irrevocably. When just one bidder is left, that bidder clinches the object at the current price. It is obvious that you should keep bidding if the price is below your true value, and should quit if the price is above. Vickrey pointed out that the second-price auction is “logically isomorphic to” the ascending auction. In the ascending auction, each bidder essentially chooses when to quit. The bidder with the highest quit-price wins, and pays the second-highest quit-price. Quit-prices in the ascending auction are equivalent to bids in the second-price auction, so it is a dominant strategy to place a bid equal to your value.
@@
Notice that Vickrey’s argument that the second-price sealed-bid auction is incentive-compatible works by first reasoning about a dynamic mechanism—the ascending auction—and then arguing that two mechanisms are equivalent. But do real participants treat these two auctions as equivalent?
@@
For auctions in the wild, it can be hard to tell whether participants are bidding truthfully, because we usually do not know their value for the object. To overcome this obstacle, Kagel et al. (1987) had lab subjects bid in auctions for a virtual prize, which had a different value for each bidder. All subjects knew their own value, in dollars, for the virtual prize. The winning bidder received cash equal to their value minus the price they paid. This method enables the experimenter to observe whether subjects are bidding truthfully.
@@
Kagel et al. (1987) randomized whether subjects participated in second-price sealed-bid auctions or in dynamic ascending auctions. Subjects played their assigned auction format 30 times in a row, with their values drawn randomly each time. Again, second-price auctions and ascending auctions are equivalent in theory; and it is a dominant strategy to choose a bid (or quit-price) equal to your value. Thus, we might expect that in both formats subjects will bid truthfully, and that the clearing price will always be equal to the second-highest value.
@@
The theoretical prediction of truthful bidding was largely borne out in ascending auctions. Subjects rapidly figured out that truthful bidding is optimal, and in 76 percent of auctions the clearing price was at the second-highest value.
@@
By contrast, in second-price sealed-bid auctions, subjects did not behave as theory predicted. Subjects often submitted bids far from their values, and the clearing price was at the second-highest value in no more than 20 percent of auctions. Even after 30 rounds of play, prices remained far from the theoretical prediction.
@@
In summary, Kagel et al. (1987) found that lab subjects rapidly converge on truthful bidding in ascending auctions, but make large and persistent mistakes in second-price auctions. This result is well-replicated (for example, see Kagel and Levin (1993); McCabe et al. (1990); Harstad (2000); Li (2017); Breitmoser and Schweighofer-Kodritsch (2022)). Ascending auctions and second-price auctions are not equivalent in practice.
@@
What did our theory miss? Why is it easy for people to see that the ascending auction is strategy-proof? Why is it hard to see this for the second-price auction?
@@
Imagine a participant who is bidding in a second-price sealed-bid auction. Let us assume that this participant is not a trained economist, and thus does not realize that the second-price auction is “logically isomorphic to” the ascending auction. [^5] In a lab experiment, Breitmoser and Schweighofer-Kodritsch (2022) found that subjects are more likely to bid truthfully in second-price auctions when the mechanism is framed in a way that makes the isomorphism salient. But why does the ascending auction framing make it easier to see the dominant strategy? Suppose that the participant values the object at \$40. To see that bidding \$40 is better than deviating to a different bid, the participant must make a case-by-case comparison, calculating payoffs for each profile of opponent bids. In particular, the participant must understand that if a truthful bid of \$40 would win at a price of \$30 (because the highest other bid is \$30), then deviating to bid \$50 would also lead to a win at \$30. And if a truthful bid of \$40 would lose, then a bid of \$50 could only win at prices above \$40. If the participant does not keep track of each contingency, then they might be tempting to bid above their true value in order to raise the chance of winning. Of course, this strategy would be a mistake, because boosting their bid to \$50 raises their chance of winning only when the highest other bid is between \$40 and \$50. In that case, the price is so high that winning is undesirable.
@@
Now imagine a participant who is bidding in a dynamic ascending auction, and values the object at \$40. The price is now \$30. If the participant continues to bid truthfully, then the participant will either win at a price between \$30 and \$40, or will lose and have payoff zero. Thus, every possible payoff from bidding truthfully is at least zero. On the other hand, if the participant deviates to quit right now, then their payoff is zero for sure. [^6] We have ignored the possibility of ties. Formally, we are modeling the ascending auction as a mechanism in which we cycle between the active bidders in some fixed order, asking each to raise their bid by a dollar or to quit irrevocably. When just one bidder remains, that bidder wins and pays their last bid. The worst-case payoff from truthful bidding is at least as good as the best-case payoff from deviating, so the participant need not keep track of different contingencies to see that truthful bidding is optimal.
@@
A symmetric argument establishes that the participant should quit when the price exceeds their value. Suppose that the participant values the object at \$40 and the price rises to \$41. If the participant quits, then their payoff is zero for sure. If instead the participant keeps bidding, then their payoff can never be positive, because now they can only win at prices above \$41.
In summary, we can distinguish dynamic ascending auctions and second-price sealed-bid auctions in this way: To play optimally in a second-price auction, you need to compare strategies case-by-case, whereas to play optimally in an ascending auction, you need not make that comparison.
@@
In Li (2017), I generalized this idea to any extensive-form mechanism. A strategy is obviously dominant if, for any deviating strategy, at any information set where the two strategies choose different actions for the first time, the worst-case payoff under the dominant strategy is at least as good as the best-case payoff under the deviation. (When the worst case from the dominant strategy is preferred to the best case from the deviation, a case-by-case comparison is not needed.) A mechanism is obviously strategy-proof if the truthful strategy is obviously dominant.
@@
Dyamic ascending auctions are obviously strategy-proof. In contrast, at the moment you submit your bid in a second-price sealed-bid auction, bidding your value can lead to a zero payoff (if you lose), whereas bidding above your value can yield a positive payoff (depending on the second-price bid). Thus, truthful bidding is not an obviously dominant strategy, and second-price auctions are not obviously strategy-proof.
@@
Obvious strategy-proofness characterizes the ascending auction. Suppose that we restrict attention to auctions that are efficient (allocating the object to a bidder with maximal value) and in which only the winning bidder makes payments. In that class, an extensive-form mechanism is obviously strategy-proof if and only if it is an ascending auction (Li and Lo, 2023, p. 576–582).
@@
Under weaker restrictions, one can characterize obviously strategy-proof mechanisms as ‘personal-clock auctions’, a larger class that includes ascending auctions, descending-price reverse auctions, and some hybrid formats (Li, 2017). Roughly, in a personal-clock auction, each participant chooses between clinching a sure thing (for example, leaving empty-handed) and a tentative alternative with a price that gets steadily worse for the participant (for example, bidding to buy the object at a price that increases each round). For another example, consider a reverse auction in which the bidder is a seller, facing a gradually descending price. At each point, the seller can either quit and keep their object (the sure thing) or offer to sell at the current price (the tentative alternative).
In a practical application of these issues, the US Federal Communications Commission auctioned off US\$19.8 billion of radio spectrum rights in 2017. It included a reverse auction that paid television stations to relinquish their over-the-air broadcast rights, so that this spectrum could be repurposed for wireless Internet access.
The Federal Communications Commission’s reverse auction had to deal with complicated feasibility constraints. To check whether an allocation is feasible, one must solve a graph coloring problem for a graph with up to 3,000 nodes and 2.7 million edges; the nodes represent television stations and the edges represent interference constraints. Stations that share an edge cannot be assigned to the same channel. This problem is computationally hard (Karp, 1972). Consequentially, the reverse auction used a suite of state-of-the-art feasibility-checking algorithms to determine the allocation and payments (Leyton-Brown et al., 2017).
@@
The reverse auction proceeded, roughly, as follows: The Federal Communications Commission makes an opening offer to each television station. Each television station can either quit or continue. If it quits, then it remains on the air, retaining its broadcast rights with zero net payment. If it continues, then either it sells its rights at the current price, or the FCC makes a lower offer and the process repeats. It is an obviously dominant strategy for each television station to bid truthfully, quitting once the price drops below its value for remaining on the air. This reverse auction belongs to the class of deferred-acceptance clock auctions studied by Milgrom and Segal (2020), which have useful implications for privacy and budget-balance.
@@
I have not described how the reverse auction determines the offers or when it decides to close a sale. Those details depend on the complicated feasibility-checking algorithms, and most bidders lacked the computational horsepower to check whether the algorithms were running as intended. But even without those details, you can see that the reverse auction is incentive-compatible. This is no accident. A mechanism is obviously strategy-proof if and only if truth-telling can be seen to be dominant from a certain kind of partial description. Such a description specifies the sequences of queries that one participant might receive, the answers that are allowed, and how queries and answers map to possible outcomes (Li, 2017, Theorem 1). Notice that this description omits how the outcome depends on unobserved opponent moves; in this sense, obviously strategy-proof mechanisms formalize the idea of a dominant strategy that can be recognized without contingent reasoning.
@@
Psychologically, it is harder to account for hypothetical contingencies, than to reason about observed events (Esponda and Vespa, 2014). Dynamic mechanisms can help participants to avoid mistakes, by resolving uncertainty about what other players have done. [^7] This insight connects to a larger literature on contingent reasoning (Martínez-Marquina et al., 2019; Cohen and Li, 2022; Esponda and Vespa, 2023), reviewed by Niederle and Vespa (2023). Glazer and Rubinstein (1996) formalize another sense in which dynamic games can be easier to understand than their static counterparts.